I need to use elimination to solve each system of equations, and don't know how to do that with these as fractions.
1/x-5 - 3/y+6 = 0
2/x+7 - 1/y-3 = 0

Answer by sehar(13)

here u hv to find x & y
from ,
1/x-5 - 3/y+6 = 0
1/x-5 = 3/y+6
y+6 = 3(x-5)
y+6= 3x-15
y=3x-21....(i)

from,
2/x+7 - 1/y-3 = 0
2/x+7 = 1/y-3
2(y-3)= x+7
2y-6 = x+7
2y= x+13
y= x+13/2....(ii)
compare (i) & (ii)
3x-21= x+13/2
6x-42=x+13
6x-x=13+42
5x=55
x=11
put x=11 in (i) or (ii)
3x-21=y
3(11)-21=y
33-21=y
y=12
so, x=11 & y=12